A chemist has 100 milliliter a of 23% acid solution and some 47% acid solution. How many milliliters of the 47% solution must be added to the 23% solution to create a 27% acid solution?
Answer by lwsshak3(11628)

A chemist has 100 milliliter a of 23% acid solution and some 47% acid solution. How many milliliters of the 47% solution must be added to the 23% solution to create a 27% acid solution?
**
let x=amt of 47% solution to be added
100+x=amt of final 27% solution
..
47%x+23%*100=27%(100+x)
.47x+23=27+.27x
.2x=4
x=20
amt of 47% solution to be added=20 ml